OpenStudy (anonymous):

simplify sqrt(5) over sqrt(180)

OpenStudy (anonymous):

1/6?

OpenStudy (anonymous):

rewrite as sqrt(5/180) and simplify

OpenStudy (anonymous):

\[\frac{\sqrt{5}}{\sqrt{180}}=\sqrt{\frac{5}{180}}=\sqrt{\frac{1}{36}}=\frac{1}{6}\]

OpenStudy (anonymous):

1/6

jhonyy9 (jhonyy9):

right wan to be +/- 1/6